OpenStudy (anonymous):

Find the domain of (sqrt(3x+12))/(x^2-36)

OpenStudy (anonymous):

[-4,6) U (6,∞)

OpenStudy (anonymous):

How did you get 4?

OpenStudy (anonymous):

This is what I get x^2-36 x^2-36=/=0 x^2 =/= 36 x =/= +- 6 (-inifinity, -6) (6,infinity)

OpenStudy (anonymous):

\[\left( -\infty , 6 \right) \upsilon \left( 6 , \right)\]

OpenStudy (anonymous):

(x-6)(x+6)=0 (3(x+4))<0 (x-6)(x+6)=0 x=6,-6 Solve for x (3(x+4))<0 x<-4 x≠6 x≠-6 x≥-4 [-4,6) U (6,∞)

OpenStudy (anonymous):

Why does the numerator have restrictions?

OpenStudy (anonymous):

i think yesterday is right

OpenStudy (phi):

you do not want to take the square root of a negative number

OpenStudy (anonymous):

oh right lol

OpenStudy (anonymous):

Ok so 3x+12 > 0 3x > -12 x > -4 Why is it not (-4,-6)(6,infinity)?

OpenStudy (phi):

that looks ok. means the same thing.

OpenStudy (anonymous):

Oops i mean >= [-4,-6) (6,inifinity) How did you get only positive 6?

OpenStudy (anonymous):

How did you get (-6,6)?

OpenStudy (phi):

check that (-4,6)(6,inf)

OpenStudy (anonymous):

Wow I have no idea what you're saying :( Why don't you include -6 in (-4,6)(6,inf)

OpenStudy (phi):

because -6 is less than -4, so it is already excluded

OpenStudy (anonymous):

Oooh that makes sense, but why (-6,6)?

OpenStudy (phi):

-6 and 6 are excluded because they would cause division by zero

OpenStudy (anonymous):

Alright nvm thanks! I figured it out. It includes -4 right?

OpenStudy (phi):

and to be totally correct you should write [-4,6)(6,inf) because -4 is allowed

OpenStudy (phi):

which is mitch's very first post up top.

OpenStudy (anonymous):

Yeah, but he didn't explain it like you did :) Sorry Mitch I can only give one medal

OpenStudy (anonymous):

np